OpenStudy (jenny.andrade):

this is annoying me whats 7^8

OpenStudy (anonymous):

5764801 is the answer when you use a calculator

OpenStudy (anonymous):

7^8 = 7*7*7*7*7*7*7*7

OpenStudy (mathmale):

Try this: Break up 7^8: (7^2)^4 = 49^4 = 49 * 49 * 49 * 49. This is just one of several ways in which you can simplify this exponentiation.